Session: Feb 2012 (session 1)
Wavelengths: 5 3.6 1.3 18 6 cm
To give feedback on this entire session, click
Entire Session Feedback
(Alastair Gunn at Mon Mar 19 15:26:25 2012) Minor failures:RECE
Our K-band receiver was not cooled for this session...
(Alastair Gunn at Mon Mar 19 15:28:28 2012) Minor failures:
A problem with our HM11 cable carrying LO up the Lovell Telescope meant there were occasional decreases in Rx power by up to 10dB. These drop outs were not noted by controllers and could have occurred at any time and remain for several hours...